ON HOLOMORPHIC CURVES IN SEMI-ABELIAN VARIETIES

JUNJIRO NOGUCHI

Abstract. The algebraic degeneracy of holomorphic curves in a semi-Abelian variety omitting a divisor is proved (Lang’s conjecture generalized to semi-Abelian varieties) by making use of the jet-projection method and the logarithmic Wronskian jet differential after Siu-Yeung. We also prove a structure theorem for the locus which contains all possible image of non-constant entire holomorphic curves in a semi-Abelian variety omitting a divisor.